# Forecasting high-impact research topics via machine learning on evolving knowledge graphs

### v2

@@ Discussion
We show how to forecast the impact of future research topics. Although we view this as a significant step towards developing truly useful AI-driven assistants, achieving this goal requires numerous further advancements. Firstly, developing methods to extract more complex information from each paper will be crucial, for instance by employing hyper-graph structures that carry more information from each paper [47, 48], which has already been demonstrated to lead to exciting results in other domains[49, 2, 50]. This might also allow for the forecast of new concepts [51, 52] and their impact. Secondly, it will be interesting to approximate impact with metrics that go beyond citations – which is a crucial topic in computational sociology and the study of the science of science [1, 2]. Additionally, introducing metrics of surprise, as discussed in [53, 54], could serve as a complementary metric to citation prediction for ranking suggestions. Finally, while the suggestion of impactful new ideas might be a key component of future AI assistants, it will be crucial to study its relation to the scientific interest of working researchers.
Methods
@@
Datasets for knowledge graph construction – To compile a list of scientific concepts in natural science, we used metadata from four major preprint servers: arXiv, bioRxiv, medRxiv, and chemRxiv. The arXiv dataset can be directly downloaded from Kaggle, while metadata from bioRxiv, medRxiv, and chemRxiv are accessible through their APIs. The full methodology and code are available on the GitHub: Impact4Cast. Our comprehensive dataset encompasses approximately 2.44 million papers, including 78,084 from arXiv’s physics.optics and quant-ph categories, which were specifically utilized for identifying domain concepts.
@@
For edge generation, we used the OpenAlex database snapshot, available for download in OpenAlex bucket stored in Amazon S3. More details can be found at the OpenAlex documentation site. The complete dataset occupies around 330 GB, expanding to approximately 1.6 TB when decompressed. Our interest was specifically in journal papers that include publication time, title, abstract, and citation information. By focusing on these criteria, we managed to reduce the dataset to a more manageable gzip-compressed size of 68 GB, comprising around 92 million scientific papers.
@@
Details on concept and edge generation – From the preprint dataset of 2.44 million papers, we analyzed each article’s title and abstract using the RAKE algorithm, enhanced with additional stopwords, to extract potential concept candidates. These candidates were stored for subsequent analysis. We filtered out concepts to retain only those with two words that appeared in nine or more articles, and those with three or more words that appeared in six or more articles. This step significantly reduced the noise from the RAKE-generated concepts, yielding a refined list of 726,439 relevant concepts. To further improve the quality of these concepts, we developed an automatic tool to identify and remove common, domain-independent errors often made by RAKE. This process, which included eliminating non-conceptual phrases, verbs, ordinal numbers, conjunctions, and adverbials, resulted in a full list of 388,825 concepts.
@@
We then specifically focused on articles within the physics.optics and quant-ph categories from arXiv to extract domain-specific concepts. By iterating our entire list of concepts to these domain-specific articles, we identified 87,741 relevant concepts. Employing our specially designed automated filtering tool for further refinement, we successfully narrowed this list down to 37,960 high-quality, domain-specific concepts.
@@
We created concept pairs, or edges, from the OpenAlex dataset, by detecting when domain-specific concepts co-occurred in paper titles or abstracts. This yielded 193,977,096 concept pairs across about 21 million papers. Each edge receives a time-stamp based on its paper’s publication date, converted to the number of days since January 1, 1990. The final full knowledge graph comprises 26,010,946 unique edges after merging multiple edges between the same concept pairs. The citation information for an edge includes the paper’s yearly citations from 2012 to 2023, alongside its total citation since publication. The OpenAlex dataset excludes yearly citations older than ten years, hence the focus on this specific ten-year time frame due to the absence of data prior to 2012. For edges formed by multiple papers, the edge weight combines the annual and total citations from all contributing papers.
@@
Training details – Our neural network consists of six fully connected layers, which include four hidden layers with 600 neurons each. The network inputs are 141 features for each pair of unconnected concepts , denoted as , where each . For instance, and represent the vertex degree of concepts and for the current year, . Detailed feature description and feature generation code are available in GitHub: Impact4Cast.
@@
In our training process for the year 2016 to predict impact in 2019, we prepared a dataset comprising approximately 689 million unconnected concept pairs. The goal was to evaluate these pairs to determine whether their 3-year citation counts would have at least citations ( is impact range). From this extensive collection, we selected all positive samples (the 3-year citation counts are at least ). An equivalent number of negative samples were then randomly chosen to match the size of the positive set. The refined dataset was subsequently divided, allocating 85% for training and 15% for testing purposes. For the evaluation dataset in 2019, which aims to predict the impact by 2022, we randomly selected 10 million unconnected pairs. Our neural network was trained using the Adam optimizer with a learning rate of and a mini-batch size of 1000. In every training batch, we randomly chose an equal number of positive and negative samples from the training set. This approach was also applied to our 2019 training process for predictions into 2022, where the trained neural network is used for future forecasting.
The full dynamic knowledge graph, alongside the data required for feature preparation and evaluation, was processed using an Intel Xeon Gold 6130 CPU with 1 TiB RAM. The network training was performed on a NVIDIA Quadro RTX 6000 GPU, with 187 GiB RAM.
@@
Individual feature’s predictive ability – In Fig. 5 (a), we observe an AUC score of 0.948 for the 2019 evaluation dataset with the neural network that uses all 141 features, trained on 2016 dataset and impact range . To explore the predictive ability of individual features, we trained separate neural networks on each feature using the same 2016 dataset, and then applied the 2019 evaluation dataset to these models. This resulted in 141 individual predictions, each from a network trained on a single feature. The features were ranked by their impact predictions, shown in the Extended Data Fig. 1.
Data availability
The data in this work are publicly available via Zenodo at https://doi.org/10.5281/zenodo.10692137 [55].
Code availability
The codes for this work can be found on the GitHub https://github.com/artificial-scientist-lab/Impact4Cast.
@@ Acknowledgements
The authors thank the organizers of OpenAlex, arXiv, bioRxiv, and medRxiv, for making scientific resources freely and readily accessible. X.G acknowledges the support from the Alexander von Humboldt Foundation.

### v3

@@ Discussion
We show how to forecast the impact of future research topics. Although we view this as a significant step towards developing truly useful AI-driven assistants, achieving this goal requires numerous further advancements. Firstly, developing methods to extract more complex information from each paper will be crucial, for instance by employing hyper-graph structures that carry more information from each paper [55, 56], which has already been demonstrated to lead to exciting results in other domains [57, 2, 58]. This might also allow for the forecast of new concepts [59, 60] and their impact. Incorporating the recent dataset [61, 62] into our research could also allow us to explore more complex data structures than those used in our paper. Secondly, it will be interesting to approximate impact with metrics that go beyond citations – which is a crucial topic in computational sociology and the study of the science of science [1, 2]. Additionally, introducing metrics of surprise, as discussed in [63, 64], could serve as a complementary metric to citation prediction for ranking suggestions. Finally, while the suggestion of impactful new ideas might be a key component of future AI assistants, it will be crucial to study its relation to the scientific interest of working researchers [65].
Materials and Methods
@@
Datasets for knowledge graph construction – To compile a list of scientific concepts in natural science, we used metadata from four major preprint servers: arXiv, bioRxiv, medRxiv, and chemRxiv. The arXiv dataset can be directly downloaded from Kaggle, while metadata from bioRxiv, medRxiv, and chemRxiv are accessible through their APIs. The full methodology and codes are available on the GitHub: Impact4Cast. Our comprehensive dataset encompasses approximately 2.44 million papers, including 78,084 from arXiv’s physics.optics and quant-ph categories, which were specifically utilized for identifying domain concepts.
@@
For edge generation, we used the OpenAlex database snapshot, available for download in OpenAlex bucket. More details can be found at the OpenAlex documentation site. The complete dataset occupies around 330 GB, expanding to approximately 1.6 TB when decompressed. Our interest was specifically in scientific journal papers that include publication time, title, abstract, and citation information. By focusing on these criteria, we managed to reduce the dataset to a more manageable gzip-compressed size of 68 GB, comprising around 92 million scientific papers. From these 92 million papers, 21 million contain at least two concepts of our final concept list and can therefore for an edge in the knowledge graph.
@@
Details on concept and edge generation – From the preprint dataset of 2.44 million papers, we analyzed each article’s title and abstract using the RAKE algorithm, enhanced with additional stopwords, to extract potential concept candidates. These candidates were stored for subsequent analysis. We filtered out concepts to retain only those with two words that appeared in nine or more articles, and those with three or more words that appeared in six or more articles. This step significantly reduced the noise from the RAKE-generated concepts, yielding a refined list of 726,439 relevant concepts. To further enhance the quality of the identified concepts, we developed a suite of automatic tools designed to identify and eliminate common, domain-independent errors often associated with RAKE. In addition, we conduct a manual review to identify and eliminate any inaccuracies in the concepts. The entire process, which included eliminating non-conceptual phrases, verbs, ordinal numbers, conjunctions, and adverbials, resulted in a full list of 368,825 concepts.
@@
We then specifically focused on articles within the physics.optics and quant-ph categories from arXiv to extract domain-specific concepts. Iterating this entire list of concepts to these domain-specific articles, we identified 87,741 relevant concepts. Employing our specially designed automated filtering tool for initial refinement and then conducting a thorough manual review to remove inaccuracies, we narrowed the list down to 37,960 high-quality, domain-specific concepts.
@@
We created concept pairs, or edges, from the OpenAlex dataset, by detecting when domain-specific concepts co-occurred in paper titles or abstracts. This yielded 193,977,096 concept pairs (including multi-edges) across about 21 million papers. Each edge receives a time-stamp based on its paper’s publication date, converted to the number of days since January 1, 1990. The final full knowledge graph comprises 26,010,946 unique edges after merging multiple edges between the same concept pairs. The citation information for an edge includes the paper’s yearly citations from 2012 to 2023, alongside its total citation since publication. The OpenAlex dataset excludes yearly citations older than ten years, hence the focus on this specific ten-year time frame due to the absence of data prior to 2012. For edges formed by multiple papers, the edge weight combines the annual and total citations from all contributing papers.
@@
Training – Our neural network consists of six fully connected layers, which include four hidden layers with 600 neurons each. The network inputs are 141 features for each unconnected concept pair , denoted as , where each . For instance, and represent the vertex degree of concepts and for the current year, . Detailed feature description and feature generation code are available in GitHub: Impact4Cast.
@@
In our training process for the year 2016 to predict impact in 2019, we prepared a dataset comprising approximately 689 million unconnected concept pairs. The goal was to evaluate these pairs to determine whether their 3-year citation counts would have at least citations ( is impact range) or not. From this extensive collection, we selected all positive samples (the 3-year citation counts are at least ). An equivalent number of negative samples were then randomly chosen to match the size of the positive set. The refined dataset was subsequently divided, allocating 85% for training and 15% for testing purposes. For the evaluation dataset in 2019, which aims to predict the impact in 2022, we randomly selected 10 million unconnected pairs. Our neural network was trained using the Adam optimizer with a learning rate of and a mini-batch size of 1000. In every training batch, we randomly chose an equal number of positive and negative samples from the training set. This approach was also applied to our 2019 training process for predictions into 2022, where the trained neural network is used for future forecasting. An example loss curve is shown in Fig.9. An example for the number of positive and negative cases of the evaluation dataset (for the experiments in Fig.4) is shown in Fig.10. In Fig.11 we explain more details of the ROC curve, specifically its relation to false positive rate (FPR) and true positive rate (TPR).
The full dynamic knowledge graph, along with the data required for feature preparation and evaluation, was processed on an Intel Xeon Gold 6130 CPU with 1 TiB of RAM. However, it is not strictly necessary to have 1 TiB of RAM for this process with the relatively small concept list of 37,960; the high memory capacity was utilized for efficiency and to handle additional operations concurrently. The final domain knowledge graph in this work occupies approximately 23.12 GB of storage. It is worth noting that knowledge graphs built from larger concept lists will require more memory, as the data size and complexity increase.
The neural network training was conducted on a standard single GPU (Nvidia Quadro RTX 6000), with each training run for different impact ranges taking approximately 1.5 hours. For benchmarking, all models – except the transformer model – were run on a standard CPU (Intel Xeon Gold 6130) with memory usage below 15 GB. Each benchmarking task took roughly one hour to complete. The transformer model, however, was run on a single GPU (Nvidia Quadro RTX 6000), taking approximately 3 hours per task.
@@
Individual feature’s predictive ability – In Fig. 4 (a), we observe an AUC score of 0.948 for the 2019 evaluation dataset with the neural network that uses all 141 features, trained on 2016 dataset and impact range . To explore the predictive ability of individual features, we trained separate neural networks on each feature using the same 2016 dataset, and then applied the 2019 evaluation dataset to these models. This resulted in 141 individual predictions, each from a network trained on a single feature. The features were ranked by their impact predictions, shown in the Fig. 12.
@@ Acknowledgements
The authors thank Burak Gurlek for interesting discussions at the start of this project, and the organizers of OpenAlex, arXiv, bioRxiv, chemRxiv, and medRxiv for making scientific resources freely and readily accessible. X.G acknowledges the support from the Alexander von Humboldt Foundation.
Author Contributions X.G. and M.K. designed research; X.G. performed research and analyzed data; and X.G. and M.K. wrote the manuscript.
Competing Interests The authors declare that they have no competing financial interests.
Data and Code availability: Data is accessible on Zenodo at https://doi.org/10.5281/zenodo.10692137 [66]. Benchmark data used in our work is also available on Zenodo at https://doi.org/10.5281/zenodo.14527306 [67]. Codes for this work are available on GitHub at https://github.com/artificial-scientist-lab/Impact4Cast.